OpenStudy (djbreezy):

what is the slope of the line passing through points (1,-5) and (4,1)? A.) 2 B.) -4/5 C.) 5/4 D.) -2

OpenStudy (texaschic101):

slope = (y2 - y1) / (x2 - x1) (1, - 5)...x1 = 1 and y1 = - 5 (4, 1)....x2 = 4 and y2 = 1 now its just a matter of subbing...pay close attention to ur signs

OpenStudy (djbreezy):

I think the answer is A but im not sure @texaschic101

OpenStudy (texaschic101):

u r correct...slope is 2